Two kilograms of Columbian coffee are mixed with 1 kilogram of Brazilian coffee to produce a mixture selling for $7.00 a kilogram. If 3 kilograms of Columbian coffee were mixed with 2 kilograms of Brazilian coffee, the mixture would sell for $6.90 a kilogram. What is the value of a kilogram of each type of cofee?

Found 2 solutions by mukhopadhyay, checkley71:
Answer by mukhopadhyay(490)

Sell Price of Columbian Coffee = $x/kg
Sell Price of Brazilian Coffee per kg = 21 - 2x (hint: 3*7=21)
3x+2(21-2x)=5(6.90)
=>42-x=34.50
=>x=7.50
Answer: Columbian sells for $7.50/kg and Brazilian for $6/kg

Answer by checkley71(8403)

2C+B=3*7.00
3C+2B=5*6.90 MULTIPLY FIRST EQUATION BY -2 WE GET
-4C-2B=-2*21.00 ADDING THIS TO THE SECOND EQUATION
3C+2B=34.50 WE GET
------------
-C=-42.00+34.50 OR C=7.50 SUBSTITUTING THIS IN THE FIRST EQUATION WE GET
2*7.50+B=21.00 OR 15.00+B=21.00 OR B=21.00-15.00 OR B=6.00
PROOF 3*7.50+12.00=34.50 OR 22.50 +12.00=34.50 34.50=34.50